Can you please help me solve: 2.9^a-4 =8.1 ?
Found 2 solutions by Alan3354, tv96:
Answer by Alan3354(69443)

2.9^a-4 =8.1
Add 4
2.9^a = 12.1
---
a*log(2.9) = log(12.1)
a = log(12.1)/log(2.9)

Answer by tv96(8)

First, do Log of 2.9 on both sides.
-this will isolate the a-4.
The equation would then be a-4=1.96
Add 4 on both sides; answer will be: 5.96